OpenStudy (anonymous):

how do you factor 4x²-9 ?

OpenStudy (anonymous):

do u know a^2 -b^2 =(a-b)(a+b)

OpenStudy (anonymous):

4x²-9 =(2x)^2 -3^2 =(2x+3)(2x-3)